# Exploring the Roles of Large Language Models in Reshaping Transportation Systems: A Survey, Framework, and Roadmap

### v1

@@ 4.2 Data analyzer
Beyond text data, vision-language models (VLMs) and multimodal language models (MLMs) are also applied to understanding traffic scenarios from heterogeneous data sources (Wen et al., 2023b; Qasemi et al., 2023; Zhou and Knoll, 2024; Lohner et al., 2024; You et al., 2024; Tong and Solmaz, 2024; Esteban et al., 2025). These models can also be integrated with external knowledge for better understanding. For example, recent studies address data scarcity through retrieval-augmented architectures. RAG-Driver (Yuan et al., 2024) integrates in-context learning with a retrieval-augmentation mechanism that dynamically retrieves expert demonstrations from a database of past driving scenarios. These retrieved examples, encoded as video control signal pairs with natural language explanations, are prefixed to the input query to guide the MLM in generating human-readable driving action descriptions and justifications. Evaluations on benchmarks like BDD-X underscore the potential of combining parametric memory with external knowledge bases for robust traffic analysis.
@@ 4.4 Summary and outlook
As information processors, LLMs are revolutionizing transportation information systems through three interrelated capabilities: semantically rich context encoding, data-driven analytical reasoning, and robust multimodal fusion. From parsing scenario synthesis tasks in CARLA (Ruan et al., 2024) to predicting crash severity through narrative analysis (Fan et al., 2024), these approaches demonstrate remarkable adaptability across data modalities and application domains. While challenges persist in hallucination mitigation (Zheng et al. (2023b)) and cross-domain generalization (Yuan et al. (2024)), current studies establish a foundation for LLM-powered transportation information processing tools that balance automation with interpretability.
@@
    Integrating external information for better response: Since the capabilities of LLMs largely depend on pretraining data, they may not have specific knowledge of complex transportation problems themselves. Retrieval-augmented generation (RAG; (Guu et al., 2020)) provides a feasible approach for incorporating external knowledge or data base for more professional responses.
@@ 5 LLMs as Knowledge Encoders
LLMs can serve as repositories of implicit and explicit world knowledge (Jiang et al., 2020; Gurnee and Tegmark, 2023), enabling structured extraction, encoding, and application of domain-specific insights for transportation problems. This capability stems from their pretraining on vast corpora that contain traffic regulations, geographical semantics, and behavioral patterns. We categorize these studies of using LLMs to encode knowledge into: (1) explicit knowledge extraction for structured reasoning and (2) latent knowledge representation through embedding spaces. These studies demonstrate how to extract useful knowledge from the large parameter space of LLMs to facilitate the solving of traffic problems.
@@ 5.1 Knowledge extractor
Specialized LLMs further enhance domain-specific knowledge retention. TransGPT (Wang et al., 2024e) introduces a specialized LLM to serve as a transportation knowledge base, including two variants: TransGPT-SM for single-modal data and TransGPT-MM for multimodal data. TransGPT-SM is fine-tuned on textual transportation datasets to address tasks like traffic analysis and recommendation generation. TransGPT-MM extends this by incorporating both textual and visual data, handling tasks such as explaining traffic phenomena and answering traffic-related questions. Similarly, TrafficSafetyGPT (Zheng et al., 2023a) grounds LLMs in safety-critical contexts through supervised fine-tuning on the TrafficSafety-2K dataset, which is a curated corpus combining government-produced guidebooks and ChatGPT-generated instruction-output pairs. This alignment enables precise identification of regulatory violations in accident reports compared to general-purpose LLMs. These two methods explicitly convert LLMs to transportation knowledge bases.
Operational knowledge synthesis is exemplified by IncidentResponseGPT (Grigorev et al., 2024), which uses LLMs to create incident response plans based on traffic incident reports and regional response guidelines. It synthesizes these guidelines into a structured form using LLMs, and then combines them with real-time incident data to generate tailored, actionable traffic incident plans. Emerging frameworks such as TARGET (Deng et al., 2023) extend this paradigm by employing LLMs to automatically generate test scenarios by encoding traffic rules into a structured domain-specific language (DSL). In the TARGET framework, the LLM parses natural language traffic rules to extract key traffic components. By transforming unstructured rules into formal scenario representations, the LLM enables automated synthesis of executable test scripts in simulators like CARLA. This approach detected over 700 violations across four autonomous driving systems, demonstrating its effectiveness in translating regulatory knowledge into actionable testing scenarios while bypassing manual DSL coding.
@@ 5.2 Knowledge representation embedder
With the advent of LLMs, an emerging approach is to directly use pretrained LLMs to extract embeddings from the text prompt (He et al., 2024; Nie et al., 2025a). Specifically, He et al. (2024) attempt to elicit the inherent geospatial knowledge from pretrained LLaMA-3 model. A structured prompt that describes basic geolocation information of the POI is first derived by retrieving OpenStreetMap. Then the prompt is fed to LLMs to obtain a continuous vector. Such vectors are demonstrated to be effective in improving spatial-temporal forecasting models, including traffic prediction. Nie et al. (2025a) further extends this approach using a linear adapter and integrates it into an expert graph neural network predictor for city-wide traffic demand estimation. Evaluations also reveal the zero-shot transferability of such embeddings.
Similar principles of the LLM-based embedding approach have recently been explored in application domains. ALT-Pilot (Omama et al., 2023) uses LLMs for autonomous navigation. Using multimodal data from LiDAR and cameras onboard, combined with vision-language models (e.g., CLIP), the system enhances vehicle localization. LLMs are used to generate the embedding of language-based landmarks in the environment, enabling open-vocabulary navigation. By encoding environmental features into a shared embedding space, the system reduces localization errors in unmapped urban areas. Language-conditioned embedding techniques further enhance scenario generation and prediction tasks. LCTGen (Tan et al., 2023) utilizes LLMs to convert textual descriptions of an expected scenario into structured scene vectors that condition a Transformer model to produce realistic traffic behaviors. This alignment between language embeddings and spatiotemporal features allows generation of critical scenarios. Similarly, Zheng et al. (2024b) integrates GPT-4V-derived embeddings into motion forecasting through Transportation Context Maps (TC-Maps). By encoding global scene semantics as weighted one-hot vectors fused via cross-attention, their Motion Transformer model achieves a 0.95% mAP gain on the Waymo dataset, demonstrating that LLM-enhanced embeddings improve trajectory prediction by contextualizing local agent behaviors within holistic scene understanding.
@@ 5.3 Summary and outlook
As knowledge encoders, LLMs transform transportation data through dual mechanisms: structuring explicit domain knowledge (e.g., VTKGs in Kuang et al. (2024), protocol templates in Grigorev et al. (2024)) and encoding latent semantics into reusable embeddings (e.g., scenario vectors of LCTGen (Tan et al., 2023), LLM2Geovec in He et al. (2024)). Although specialized models such as TransGPT (Wang et al., 2024e) demonstrate the value of domain adaptation, challenges persist to maintain the freshness of knowledge. For example, incident response plans can become outdated as regulations evolve, and navigation embeddings like in ALT-Pilot (Omama et al., 2023) require continuous map updates. Nevertheless, the fusion of parametric knowledge (learned during pretraining or fine-tuning) and nonparametric representations suggests a path toward applying LLMs that dynamically internalize transportation knowledge while remaining grounded in real-world traffic problems.
@@
    Improving representation quality through alignment tuning: One of the limitations of using frozen pretrained LLMs to encoding knowledge is that the representation embedding is fixed and cannot be adaptively refined to tasks. A possible solution is to use supervised fine-tuning (Wang et al., 2023f) to improve the quality and specificity of the representations and further align them with human preference.
@@ 6.1 Function designer
Benefiting from pertained tasks such autoregressive language modeling in Eq. 2 and the in-context learning ability from textual demonstration, LLMs specialize in the design or refinement of code- or rule-based functions for traffic management. Such functions like the reward function in reinforcement learning (RL) are difficult to design manually and require expert guidance. LLMs revolutionize algorithmic design by translating natural language specifications into executable functions or codes.
@@ 6.2 World simulator
The integration of LLMs with physics engines (world models) yields unprecedented scenario customization. For example, DriveDreamer-2 (Zhao et al., 2024) employs an LLM interface to convert user queries into agent trajectories, which are then used to generate high-definition maps adhering to traffic regulations. The versatility of LLMs enable the world model to generate customized driving videos from user’s textual prompt, including uncommon scenarios like vehicles abruptly cutting in. DriveMM (Huang et al., 2024a) integrates LLMs into world models by developing a large multimodal model that synthesizes heterogeneous inputs to simulate dynamic driving environments and generate actionable outputs. It demonstrates that combining LLMs with multimodal data and structured training pipelines can produce world simulators for context-aware driving.
@@ 6.5 Summary and outlook
As generators, LLMs transform transportation systems through three key capabilities: 1) translating abstract requirements into functional algorithms (Xia et al., 2024; Han et al., 2024), 2) synchronizing photorealistic environments with controllable dynamics (Zhao et al., 2024; Gao et al., 2023b), and 3) providing interpretable system evaluations for self-refinement (Tian et al., 2024; Pang et al., 2024a). Persistent challenges include maintaining physical fidelity in the generated results. For example, while the generative world models are impressive in simulating traffic scenarios, they still trail real-world data distributions. However, the integration of LLMs with simulation engines and self-refinement mechanisms points to a future where ITS can self-generate their training ecosystems while maintaining alignment with physical and regulatory constraints.
@@ 7.1 Decision maker
In travel behavior modeling, LLMs bypass the traditional data-driven parameter calibration parameter calibration by leveraging semantic reasoning. Mo et al. (2023) demonstrates that structured prompts enable GPT-family models to match supervised models like multinomial logit models and neural networks in mode choice accuracy. However, the study notes occasional "hallucinations" in explanations where outputs lack logical consistency. Liu et al. (2024e) further evaluate the ability of LLMs to simulate human decision making in mode choice using a stated preference dataset. They first test zero-shot LLMs, finding significant behavioral misalignment due to discrepancies between LLM reasoning and real traveler preferences. To address this, they introduce persona-based few-shot learning, effectively bridging the gap between LLM reasoning patterns and empirical traveler preferences. Beyond discrete choices, Tang et al. (2024b) employs LLMs for personalized itinerary generation. LLMs generate human-readable itineraries by integrating optimized POI sequences and contextual descriptions. This hybrid method overcomes the limitations of spatial unawareness and static knowledge of standalone LLMs, ensuring both personalization and geographic feasibility. The LLM-based decision making process can also be integrated with behavior theory. In (Chen et al., 2025), the authors synthesize behavioral theory (Protective Action Decision Model), contextual cues, and a memory-based RL to enhance wildfire evacuation decision prediction. LLMs are prompted to simulate human cognitive processes, structured into threat assessment and risk perception stages. The memory module refines decisions by storing and retrieving past errors and self-reflections.
@@
Operational traffic optimization also benefits from multi-agent coordination capabilities of LLMs. CoMAL (Yao et al., 2024) integrates multiple LLM agents to tackle mixed-autonomy traffic problems by optimizing traffic flow. It employs a collaboration module where autonomous vehicles communicate using LLMs to allocate roles and discuss strategies in real-time. It demonstrates superior performance in optimizing mixed-autonomy traffic compared to RL-based models. Meanwhile, Orfanoudakis et al. (2025) combines the Decision Transformer (DT) (Chen et al., 2021) with GNNs to optimize the charging schedules of electric vehicles. GPT-2-based DT is trained to predict actions by modeling sequences of states, actions, and rewards. This hybrid approach outperforms the heuristic baselines and RL methods in the EV2Gym simulator. These works illustrate LLMs’ roles as both collaborative planners and sequential action learners in infrastructure-scale optimization.
@@ 7.3 Spatial-temporal predictor
As Transformers naturally excel at handling sequential data, LLMs have shown outstanding performance in time series analysis (Jin et al., 2023a; Gruver et al., 2023) as well as in spatial-temporal data mining (Jin et al., 2023b; Li et al., 2024c). One of the most fundamental macroscopic quantities in transportation systems is the traffic flow, usually structured as spatial-temporal data. Traffic flow forecasting is one of the fundamental tasks in data-centric transportation systems. Traditional methods include using convolutional neural networks (CNNs), recurrent neural networks (RNNs), and graph neural networks (GNNs) to model graph-based traffic time series (Tedjopurnomo et al., 2020; Yin et al., 2021; Xue et al., 2025). MLP-based (Shao et al., 2022; Qin et al., 2023; Nie et al., 2025b) and Transformer-based (Xu et al., 2020; Yan et al., 2021; Liu et al., 2023a) architectures have emerged as new alternatives.
@@
More recently, GPT-like architectures have been adopted as backbone forecasters for traffic flow. ST-LLM (Liu et al., 2024b) introduces an embedding module to create a unified spatial-temporal representation and feeds the embeddings to LLMs to predict future traffic time series. To adapt the LLM to traffic prediction tasks, ST-LLM employs a partially frozen attention mechanism, where the frozen layers preserve foundational knowledge, and unfrozen attention layers focus on capturing the specific dependencies. To bridge the gap between sequential text and traffic data, STG-LLM (Liu et al., 2024d) introduces a spatial-temporal graph tokenizer that transforms traffic data into tokens. This transformation reduces the complexity of the graph-structured data, making it more accessible for LLMs. For explainability, xTP-LLM (Guo et al., 2024) transforms multimodal traffic information into natural language descriptions. CoT prompts are used to guide LLMs in identifying relevant factors from the given information. Then LLaMA-2 is fine-tuned using language-based instructions to align with the specific requirements of traffic prediction. Empirical evaluations show that xTP-LLM not only achieves competitive accuracy, but also provides intuitive explanations for its predictions. Both of the above methods utilize the fine-tuning technique to adapt LLMs to the traffic domain. To improve efficiency and reduce computational demands, TPLLM (Ren et al., 2024) further introduces a low-rank adaptation (LoRA) fine-tuning approach for GPT-2, allowing effective learning with fewer parameters. These architectures are also applied to traffic data imputation tasks (Chen et al., 2023c; Zhang et al., 2024c; Nie et al., 2024a; Fang et al., 2025).
@@
Moreover, LLMs have also been used as backbone predictors for trajectory prediction by decoding mobility patterns across scales (Wang et al., 2023d; Xue et al., 2024; Liang et al., 2024a; Haydari et al., 2024; Long et al., 2024b; Zhu et al., 2024; Zhang et al., 2024g). This can be categorized as vehicle trajectory modeling from a micro-perspective and mobility prediction from a macro-perspective. For microscopic trajectory modeling, existing studies explore different trajectory encoding strategies for LLMs. Specifically, LMTraj (Bae et al., 2024) transforms trajectory prediction into a QA problem. They convert pedestrian trajectory coordinates and scene images into textual prompts using numerical tokenizers, and integrate them into a QA template for LLMs. Chib and Singh (2024) presents LG-Traj, a method that uses an LLM-based architecture with a motion encoder to capture motion patterns and a social decoder to capture social interactions among pedestrians. In (Lan et al., 2024), Traj-LLM is proposed to leverage pretrained LLMs without explicit prompt engineering. The approach begins with sparse context joint encoding to process agent and scene features into a form comprehensible by LLMs. In addition, LC-LLM (Peng et al., 2024) reformulates lane change prediction as a language modeling problem. It processes heterogeneous driving scenario information in natural language prompts for LLMs and employs supervised fine-tuning to tailor the LLM for lane change prediction tasks.

### v2

@@ 4.2 Data analyzer
Beyond text data, vision-language models (VLMs) and multimodal language models (MLMs) are also applied to understanding traffic scenarios from heterogeneous data sources (Wen et al., 2023b; Qasemi et al., 2023; Zhou and Knoll, 2024; Lohner et al., 2024; You et al., 2024; Tong and Solmaz, 2024; Esteban et al., 2025). These models can also be integrated with external knowledge for better understanding. For example, recent studies address data scarcity through retrieval-augmented architectures. RAG-Driver (Yuan et al., 2024) integrates ICL with a retrieval-augmentation mechanism that dynamically retrieves expert demonstrations from a database of past driving scenarios. These retrieved examples, encoded as video control signal pairs with natural language explanations, are prefixed to the input query to guide the MLM in generating human-readable driving action descriptions and justifications. Evaluations on benchmarks like BDD-X underscore the potential of combining parametric memory with external knowledge bases for robust traffic analysis.
@@ 4.4 Summary and outlook
As information processors, LLMs are revolutionizing transportation information systems through three interrelated capabilities: semantically rich context encoding, data-driven analytical reasoning, and robust multimodal fusion. From parsing scenario synthesis tasks in CARLA (Ruan et al., 2024) to predicting crash severity through narrative analysis (Fan et al., 2024), these approaches demonstrate remarkable adaptability across data modalities and application domains. In summary, current studies establish a foundation for LLM-powered transportation information processing tools that balance automation with interpretability. However, several critical limitations persist in current research that constrain the broader applicability of them. First, most existing implementations rely heavily on benchmark datasets or constrained simulation settings, limiting their robustness in real-world, noisy, or out-of-distribution environments. For instance, context encoders often assume grammatically correct and unambiguous user input, which may not reflect field deployment conditions. Furthermore, multimodal fusion approaches typically focus on aligning no more than two data modalities and lack comprehensive evaluations on diverse and real-time traffic data streams. Second, many proposed models exhibit a strong dependence on prompt engineering or fine-tuning strategies, leading to reproducibility and generalization concerns. The lack of standardized evaluation protocols for transportation-specific tasks exacerbates this issue. Despite the progress of frameworks like ChatSUMO and SeeUnsafe, their underlying LLMs are still vulnerable to hallucinations and fail to guarantee interpretability in high-stakes scenarios such as accident analysis or transport planning.
@@
    Integrating external information for better response: Since the capabilities of LLMs largely depend on pretraining data, they may not have specific knowledge of complex transportation problems themselves. RAG (Guu et al., 2020) provides a feasible approach for incorporating external knowledge or data base for more professional responses.
  3. 3.
    Robust multimodal aligning in real-world scenarios: Future work should investigate LLM-based architectures capable of operating under noisy, incomplete, or ambiguous real-world inputs. This may require robust data alignment, continual learning, and probabilistic inference mechanisms that can respond to dynamic and uncertain environments.
@@ 5 LLMs as Knowledge Encoders
LLMs can serve as repositories of implicit and explicit world knowledge (Jiang et al., 2020; Gurnee and Tegmark, 2023), enabling structured extraction, encoding, and application of domain-specific insights for transportation problems. This capability stems from their pretraining on vast corpora that contain traffic regulations, geographical semantics, and behavioral patterns. Fig. 5 categorizes these studies of using LLMs to encode knowledge into: (1) explicit knowledge extraction for structured reasoning and (2) latent knowledge representation through embedding spaces. These studies demonstrate how to extract useful knowledge from the large parameter space of LLMs to facilitate the solving of traffic problems.
@@ 5.1 Knowledge extractor
Specialized LLMs further enhance domain-specific knowledge retention. TransGPT (Wang et al., 2024e) introduces a specialized LLM to serve as a transportation knowledge base, including two variants: TransGPT-SM for single-modal data and TransGPT-MM for multimodal data. TransGPT-SM is fine-tuned on textual transportation datasets to address tasks like traffic analysis and recommendation generation. TransGPT-MM extends this by incorporating both textual and visual data, handling tasks such as explaining traffic phenomena and answering traffic-related questions. Similarly, TrafficSafetyGPT (Zheng et al., 2023a) grounds LLMs in safety-critical contexts through SFT on the TrafficSafety-2K dataset, which is a curated corpus combining government-produced guidebooks and ChatGPT-generated instruction-output pairs. This alignment enables precise identification of regulatory violations in accident reports compared to general-purpose LLMs. These two methods explicitly convert LLMs to transportation knowledge bases.
Operational knowledge synthesis is exemplified by IncidentResponseGPT (Grigorev et al., 2024), which uses LLMs to create incident response plans based on traffic incident reports and regional response guidelines. It synthesizes these guidelines into a structured form using LLMs, and then combines them with real-time incident data to generate tailored, actionable traffic incident plans. Emerging frameworks such as TARGET (Deng et al., 2023) extend this paradigm by employing LLMs to automatically generate test scenarios by encoding traffic rules into a structured domain-specific language. In the TARGET framework, the LLM parses natural language traffic rules to extract key traffic components. By transforming unstructured rules into formal scenario representations, the LLM enables automated synthesis of executable test scripts in simulators like CARLA. This approach detected over 700 violations across four autonomous driving systems, demonstrating its effectiveness in translating regulatory knowledge into actionable testing scenarios while bypassing manual DSL coding.
@@ 5.2 Knowledge representation embedder
With the advent of LLMs, an emerging approach is to directly use pretrained LLMs to extract embeddings from the text prompt (He et al., 2024; Nie et al., 2025a). Specifically, He et al. (2024) attempt to elicit the inherent geospatial knowledge from pretrained LLaMA-3 model. A structured prompt that describes basic geolocation information of the POI is first derived by retrieving OpenStreetMap. Then the prompt is fed to LLMs to obtain a continuous vector. Such vectors are demonstrated to be effective in improving spatial-temporal forecasting models, including traffic prediction. Nie et al. (2025a) further extends this approach using a linear adapter and integrates it into an expert GNN predictor for city-wide traffic demand estimation. Evaluations also reveal the zero-shot transferability of such embeddings.
Similar principles of the LLM-based embedding approach have recently been explored in application domains. ALT-Pilot (Omama et al., 2023) uses LLMs for autonomous navigation. Using multimodal data from LiDAR and cameras onboard, combined with VLMs (e.g., CLIP), the system enhances vehicle localization. LLMs are used to generate the embedding of language-based landmarks in the environment, enabling open-vocabulary navigation. By encoding environmental features into a shared embedding space, the system reduces localization errors in unmapped urban areas. Language-conditioned embedding techniques further enhance scenario generation and prediction tasks. LCTGen (Tan et al., 2023) utilizes LLMs to convert textual descriptions of an expected scenario into structured scene vectors that condition a Transformer model to produce realistic traffic behaviors. This alignment between language embeddings and spatiotemporal features allows generation of critical scenarios. Similarly, Zheng et al. (2024b) integrates GPT-4V-derived embeddings into motion forecasting through Transportation Context Maps. By encoding global scene semantics as weighted one-hot vectors fused via cross-attention, their Motion Transformer model achieves a 0.95% mAP gain on the Waymo dataset, demonstrating that LLM-enhanced embeddings improve trajectory prediction by contextualizing local agent behaviors within holistic scene understanding.
@@ 5.3 Summary and outlook
As knowledge encoders, LLMs transform transportation data through dual mechanisms: structuring explicit domain knowledge (e.g., VTKGs in Kuang et al. (2024), protocol templates in Grigorev et al. (2024)) and encoding latent semantics into reusable embeddings (e.g., scenario vectors of LCTGen (Tan et al., 2023), LLM2Geovec in He et al. (2024)). Although specialized models such as TransGPT (Wang et al., 2024e) demonstrate the value of domain adaptation, challenges persist to maintain the freshness of knowledge. For example, incident response plans can become outdated as regulations evolve, and navigation embeddings like in ALT-Pilot (Omama et al., 2023) require continuous map updates. Nevertheless, the fusion of parametric knowledge (learned during pretraining or fine-tuning) and nonparametric representations suggests a path toward applying LLMs that dynamically internalize transportation knowledge while remaining grounded in real-world traffic problems. However, key limitations remain unaddressed. First, static fine-tuning or prompt-based extraction methods cannot keep up with evolving traffic regulations, infrastructure changes, or emerging mobility patterns. For example, an incident plan generated by IncidentResponseGPT can quickly become outdated when local response guidelines are updated. There is no standard mechanism to incrementally update embeddings or graphs as new data arrives. Second, there is no unified benchmark or tool to measure knowledge quality (e.g., correctness, coverage) or embedding utility across different transportation tasks. Finally, interpretability of learned representations is also underexplored, hindering trust and adoption by domain experts.
@@
    Improving representation quality through alignment tuning: One of the limitations of using frozen pretrained LLMs to encoding knowledge is that the representation embedding is fixed and cannot be adaptively refined to tasks. A possible solution is to use SFT (Wang et al., 2023f) to improve the quality and specificity of the representations and further align them with human preference.
@@ 6.1 Function designer
Benefiting from pertained tasks such autoregressive language modeling in Eq. 2 and the ICL ability from textual demonstration, LLMs specialize in the design or refinement of code- or rule-based functions for traffic management. Such functions like the reward function in RL are difficult to design manually and require expert guidance. LLMs revolutionize algorithmic design by translating natural language specifications into executable functions or codes.
@@ 6.2 World simulator
The integration of LLMs with physics engines (world models) yields unprecedented scenario customization. For example, DriveDreamer-2 (Zhao et al., 2024a) employs an LLM interface to convert user queries into agent trajectories, which are then used to generate high-definition maps adhering to traffic regulations. The versatility of LLMs enable the world model to generate customized driving videos from user’s textual prompt, including uncommon scenarios like vehicles abruptly cutting in. DriveMM (Huang et al., 2024a) integrates LLMs into world models by developing a large multimodal model that synthesizes heterogeneous inputs to simulate dynamic driving environments and generate actionable outputs. It demonstrates that combining LLMs with multimodal data and structured training pipelines can produce world simulators for context-aware driving.
@@ 6.5 Summary and outlook
As generators, LLMs transform transportation systems through three key capabilities: 1) translating abstract requirements into functional algorithms (Xia et al., 2024; Han et al., 2024), 2) synchronizing photorealistic environments with controllable dynamics (Zhao et al., 2024a; Gao et al., 2023b), and 3) providing interpretable system evaluations for self-refinement (Tian et al., 2024; Pang et al., 2024a). Maintaining physical fidelity in the generated results is crucial for real-world deployment. For example, while the generative world models are impressive in simulating traffic scenarios, they still trail real-world data distributions. However, the integration of LLMs with simulation engines and self-refinement mechanisms points to a future where ITS can self-generate their training ecosystems while maintaining alignment with physical and regulatory constraints. Despite these advances, important limitations constrain their deployment. First, generated functions and synthetic environments often diverge from real-world physics or omit critical semantics. For example, world models may misrepresent vehicle dynamics under extreme maneuvers, and synthetic scenarios can lack corner-case diversity beyond prompt scope. Second, high-quality generation, especially photorealistic video or complex code, can incur prohibitive inference costs and latency, limiting real-time or embedded use. Third, without rigorous validation, the generated components can encode unsafe behaviors or reflect biases from training data, posing risks in safety-critical contexts.
@@
  4. 4.
    Physics-aware generation pipelines. Incorporate differentiable physics engines or domain constraints into the generation loop, e.g., embedding vehicle dynamics simulators into LLM prompting, to ensure physically plausible code and components.
@@ 7.1 Decision maker
In travel behavior modeling, LLMs bypass the traditional data-driven parameter calibration parameter calibration by leveraging semantic reasoning. Mo et al. (2023) demonstrates that structured prompts enable GPT-family models to match supervised models like multinomial logit models and neural networks in mode choice accuracy. However, the study notes occasional "hallucinations" in explanations where outputs lack logical consistency. Liu et al. (2024f) further evaluate the ability of LLMs to simulate human decision making in mode choice using a stated preference dataset. They first test zero-shot LLMs, finding significant behavioral misalignment due to discrepancies between LLM reasoning and real traveler preferences. To address this, they introduce persona-based few-shot learning, effectively bridging the gap between LLM reasoning patterns and empirical traveler preferences. Beyond discrete choices, Tang et al. (2024b) employs LLMs for personalized itinerary generation. LLMs generate human-readable itineraries by integrating optimized POI sequences and contextual descriptions. This hybrid method overcomes the limitations of spatial unawareness and static knowledge of standalone LLMs, ensuring both personalization and geographic feasibility. The LLM-based decision making process can also be integrated with behavior theory. In (Chen et al., 2025), the authors synthesize behavioral theory (Protective Action Decision Model), contextual cues, and a memory-based RL to enhance wildfire evacuation decision prediction. LLMs are prompted to simulate human cognitive processes, structured into threat assessment and risk perception stages. The memory module refines decisions by storing and retrieving past errors and self-reflections.
@@
Operational traffic optimization also benefits from multi-agent coordination capabilities of LLMs. CoMAL (Yao et al., 2024a) integrates multiple LLM agents to tackle mixed-autonomy traffic problems by optimizing traffic flow. It employs a collaboration module where autonomous vehicles communicate using LLMs to allocate roles and discuss strategies in real-time. It demonstrates superior performance in optimizing mixed-autonomy traffic compared to RL-based models. Meanwhile, Orfanoudakis et al. (2025) combines the Decision Transformer (DT) (Chen et al., 2021) with GNNs to optimize the charging schedules of electric vehicles. GPT-2-based DT is trained to predict actions by modeling sequences of states, actions, and rewards. This hybrid approach outperforms the heuristic baselines and RL methods in the EV2Gym simulator. These works illustrate LLMs’ roles as both collaborative planners and sequential action learners in infrastructure-scale optimization.
@@ 7.3 Spatial-temporal predictor
As Transformers naturally excel at handling sequential data, LLMs have shown outstanding performance in time series analysis (Jin et al., 2023a; Gruver et al., 2023) as well as in spatial-temporal data mining (Jin et al., 2023b; Li et al., 2024c). One of the most fundamental macroscopic quantities in transportation systems is the traffic flow, usually structured as spatial-temporal data. Traffic flow forecasting is one of the fundamental tasks in data-centric transportation systems. Traditional methods include using convolutional neural networks (CNNs), recurrent neural networks (RNNs), and GNNs to model graph-based traffic time series (Tedjopurnomo et al., 2020; Yin et al., 2021; Xue et al., 2025). MLP-based (Shao et al., 2022; Qin et al., 2023; Nie et al., 2025b) and Transformer-based (Xu et al., 2020; Yan et al., 2021; Liu et al., 2023a) architectures have emerged as new alternatives.
@@
More recently, GPT-like architectures have been adopted as backbone forecasters for traffic flow. ST-LLM (Liu et al., 2024c) introduces an embedding module to create a unified spatial-temporal representation and feeds the embeddings to LLMs to predict future traffic time series. To adapt the LLM to traffic prediction tasks, ST-LLM employs a partially frozen attention mechanism, where the frozen layers preserve foundational knowledge, and unfrozen attention layers focus on capturing the specific dependencies. To bridge the gap between sequential text and traffic data, STG-LLM (Liu et al., 2024e) introduces a spatial-temporal graph tokenizer that transforms traffic data into tokens. This transformation reduces the complexity of the graph-structured data, making it more accessible for LLMs. For explainability, xTP-LLM (Guo et al., 2024) transforms multimodal traffic information into natural language descriptions. CoT prompts are used to guide LLMs in identifying relevant factors from the given information. Then LLaMA-2 is fine-tuned using language-based instructions to align with the specific requirements of traffic prediction. Empirical evaluations show that xTP-LLM not only achieves competitive accuracy, but also provides intuitive explanations for its predictions. Both of the above methods utilize the fine-tuning technique to adapt LLMs to the traffic domain. To improve efficiency and reduce computational demands, TPLLM (Ren et al., 2024) further introduces a LoRA fine-tuning approach for GPT-2, allowing effective learning with fewer parameters. These architectures are also applied to traffic data imputation tasks (Chen et al., 2023c; Zhang et al., 2024c; Nie et al., 2024a; Fang et al., 2025).
@@
Moreover, LLMs have also been used as backbone predictors for trajectory prediction by decoding mobility patterns across scales (Wang et al., 2023d; Xue et al., 2024; Liang et al., 2024a; Haydari et al., 2024; Long et al., 2024b; Zhu et al., 2024; Zhang et al., 2024g). This can be categorized as vehicle trajectory modeling from a micro-perspective and mobility prediction from a macro-perspective. For microscopic trajectory modeling, existing studies explore different trajectory encoding strategies for LLMs. Specifically, LMTraj (Bae et al., 2024) transforms trajectory prediction into a QA problem. They convert pedestrian trajectory coordinates and scene images into textual prompts using numerical tokenizers, and integrate them into a QA template for LLMs. Chib and Singh (2024) presents LG-Traj, a method that uses an LLM-based architecture with a motion encoder to capture motion patterns and a social decoder to capture social interactions among pedestrians. In (Lan et al., 2024), Traj-LLM is proposed to leverage pretrained LLMs without explicit prompt engineering. The approach begins with sparse context joint encoding to process agent and scene features into a form comprehensible by LLMs. In addition, LC-LLM (Peng et al., 2024) reformulates lane change prediction as a language modeling problem. It processes heterogeneous driving scenario information in natural language prompts for LLMs and employs SFT to tailor the LLM for lane change prediction tasks.